Suppose a boil water notice is sent out advising all residents in the area to boil their water before drinking it or using it for cooking. You need to boil 11.5 L

11.5 L of water using your natural gas (primarily methane) stove. What volume of natural gas is needed to boil the water if only 17.1%

17.1% of the heat generated goes towards heating the water. Assume the density of methane is 0.668 g/L

0.668 g/L, the density of water is 1.00 g/mL

1.00 g/mL, and that the water has an initial temperature of 21.0 °C

21.0 °C. Enthalpy of formation values can be found in this table. Assume that gaseous water is formed in the combustion of methane.


1
Expert's answer
2021-03-15T09:26:35-0400

Energy required for boiling of "11.5 L" of water:

"1. \" Energy required to increase it's temperature from "21\degree C\to100\degree C"

"2. \" Energy required for boiling(Latent heat of evaporation)

"=mC(T_2-T_1)+mL_v=" "11.5 \times 4.184\times (100-21)+11.5\times 2256=" "29745.16J"

Let "V" be the volume of methane required,

Mass of methane will be"=d\times V=0.668V g" (where d is the density)

Calorific value of methane"=55KJ\/g"

"17.1" % of Total energy of methane "=" "29745.16J"

"\implies \frac{17.1}{100}\times 55000\times 0.668\times V=29745.16J"

"V=4.73L"

"V=4.73L"